Show that the explicit sequence {yn}∞n=0 such that yn = A(2n )+ B(-1)n for any nonzero

constants A and B is the solution of the recurrence relation

yn = yn-1 + 2yn-2 for n >1.


1
Expert's answer
2022-02-22T11:29:58-0500

Let us solve the reccurence relation

"y_n = y_{n-1 }+ 2y_{n-2}."

Its characteristic equation "k^2=k+2" is equivalent to "k^2-k-2=0," and hence to "(k-2)(k+1)=0."

Therefore, the last equation has the roots "k_1=2,\ k_2=-1."

We conclude that the general solution of the reccurence relation is of the form:

"y_n=A\cdot 2^n+B\cdot(-1)^n."